OpenStudy (anonymous):

Answer gets a medal (pic included) What is tan A? A. 4/5 B. 3/5 C. 4/3 D. 3/4

OpenStudy (anonymous):



OpenStudy (anonymous):

So Tan A = BC/AC u can find it accordingly; nt allowed to give direct solutions here

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

we are NOT allowed to give the direct spoonfeed answer..we can guide the step but you have to do calculations

OpenStudy (anonymous):

how do I figure it out though?

OpenStudy (anonymous):

well Tan of any angle is the length of opposite side divided by length of adjacent side. so u know which is opposite and which is adjacent side?

OpenStudy (anonymous):

is 8 adjacent and 6 opposite

OpenStudy (anonymous):

if so then we have 8/10

OpenStudy (anonymous):

actually that is the mistake people do. that is for angle B. but for angle A, the adjacent side is AC which is 6, and opposite is BC which is 8. see from the angle A ..8 units wont look like adjacent

OpenStudy (anonymous):

so 6/10 ?

OpenStudy (anonymous):

well 10 is the hypotenuse we don't touch that. the adjacent side is NEVER the hypotenuse. so Tan A = BC/AC which is 8/6.

OpenStudy (anonymous):

and we simplify to 4/3 correct?

OpenStudy (anonymous):

which is answer C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So for others such as cos, sin and tan NEVER touch hypotenuse

OpenStudy (anonymous):

cos is base/hypotenuse so u consider it. sin is opposite/hypotenuse. but Tan = opposite/adjacent and cot = 1/Tan in these 2 cases u dnt take it

OpenStudy (anonymous):

what would ok lets say I have  How would i do this